OpenStudy (anonymous):

For two quadrilaterals A B C D is similar to Q R S T, line segment A B equals 5 inches, line segment B C equals line segment C D equals 8 inches, and line segment D A equals 2 inches. Match the given scale factor with the given side length between quadrilaterals A B C D and Q R S T. Match Term Definition 1.25 A) line segment S T equals 4 inches 3 B) line segment Q R equals ten inches 2 C) line segment R S equals ten inches 0.5 D) line segment T Q equals 6 inches

OpenStudy (babalooo):

okay I GOT THIS

OpenStudy (anonymous):

lol kk

OpenStudy (babalooo):



OpenStudy (babalooo):

Okay so i'm using one quad for both properties, with ABCD info on the outside and the QRST on the inside

OpenStudy (anonymous):

ok lol but im still confused

OpenStudy (babalooo):

The only way to get the scale factor to match the given values was to divide the bigger value by the smaller value

OpenStudy (babalooo):

oops sorry

OpenStudy (babalooo):

I meant the the QRST length/ABCD length

OpenStudy (babalooo):

So like for a simple example, scale factor is a multiple of a certain value. Say, A=3 and B=6, we know b is 2 times A value, so the scale factor is 2. 6/3=2. It is the same in this case. If you A and B in the simple example I proposed, as a length (say A=ST and B=AB) we have A=4 and B=8 correct? to find the scale factor, we do A/B=4/8=0.5

OpenStudy (babalooo):

Apply method to find scale factor of all sides. Hope this makes sense

OpenStudy (anonymous):

ok thanks @babalooo

OpenStudy (babalooo):

np